OpenStudy (anonymous):

givin the fuction if f(x)=3x-4/5 which of the below expressions are correct? f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Okay so this in inverse functions is this \[\large f(x) = 3x - \frac{4}{5}\] or \[\large f(x) = \frac{3x - 4}{5}\]

OpenStudy (anonymous):

okayy

OpenStudy (johnweldon1993):

There was a question in that :P

OpenStudy (anonymous):

ohh my bad lol its the bottom one sorry

OpenStudy (johnweldon1993):

Lol alright so we have \[\large y = \frac{3x - 4}{5}\] Im sure you know that f(x) can just be written as 'y' right?

OpenStudy (anonymous):

yes sir :)

OpenStudy (johnweldon1993):

Good..so yes we have \[\large y = \frac{3x - 4}{5}\] what the basic idea of inverse function is...you switch the 'x' ad the 'y'....then you solve again for 'y' so...lets switch x and y \[\large x = \frac{3y - 4}{5}\]

OpenStudy (johnweldon1993):

And now we just need to use algebra to solve this for 'y' again...how would we do that?

OpenStudy (anonymous):

multiply the x by 5 nd 3y-4 by 5?

OpenStudy (johnweldon1993):

Correct....leaving us with \[\large 5x = 3y - 4\] correct?

OpenStudy (anonymous):

yeppers

OpenStudy (johnweldon1993):

What next? we need 'y' all by itself

OpenStudy (anonymous):

divide all of the numbers by 3 or is it just 5/3

OpenStudy (johnweldon1993):

Hmm...well yeah we could do that first....but maybe we should add 4 to both sides first ...what do ya think?

OpenStudy (anonymous):

suree which ever way is easier for me lol

OpenStudy (anonymous):

you would get 8x = 3y right ?

OpenStudy (johnweldon1993):

Well we COULD do it either way...lol but now you've forced me to show you both ways :P

OpenStudy (anonymous):

nahh just do anyone idc and i didnt force you :p

OpenStudy (johnweldon1993):

Lol :P alright well we'll do it the other way I said \[\large 5x = 3y - 4\] after you add 4 to both sides you get \[\large 5x + 4 = 3y\] right?

OpenStudy (anonymous):

yaa

OpenStudy (johnweldon1993):

And finally you would divide everything by 3 to solve for 'y' \[\large f^{-1}(x) = \frac{5x + 4}{3}\]

OpenStudy (anonymous):

Omg i love you forever jk u are amazing

OpenStudy (johnweldon1993):

Lol <3 haha glad to help :)

OpenStudy (anonymous):

can i tag u if i have anymore questions please u make me feel not stupid

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

thankss